# Content-Oblivious Leader Election in 2-Edge-Connected Networks

### v2

@@ Abstract
In this work, we present an asynchronous content-oblivious leader election algorithm that quiescently terminates in any 2-edge connected network with message complexity $O(m\cdot N\cdot\mathsf{ID_{\min}})$, where $m$ is the number of edges, $N$ is a known upper bound on the number of nodes, and $\mathsf{ID_{\min}}$ is the smallest $\mathsf{ID}$. Combined with the previous simulation result, our finding implies that any algorithm from the noiseless setting can be simulated in the fully-defective setting without assuming a preselected leader, entirely refuting the original conjecture.
@@ 1 Introduction
Fault tolerance is a cornerstone of distributed computing, enabling systems to remain operational despite various failures, such as node crashes or channel noise [Barborak1993, Dubrova2013, Koren2020, Raynal2018]. One important category of faults, known as alteration errors, changes the content of messages but does not create new messages or eliminate existing ones. Traditional methods for handling such errors often rely on adding redundancy through coding techniques and assume known bounds on the frequency or severity of faults. However, in many practical settings, such assumptions may not hold, limiting the effectiveness of these approaches.
@@
Censor-Hillel, Cohen, Gelles, and Sela [fully-defective-22] introduced the model of fully-defective asynchronous networks, in which all links are subject to a severe form of alteration errors: The content of every message can be arbitrarily modified, although the adversary cannot insert new messages or remove existing messages. Since messages can no longer carry meaningful information, the model is equivalent to content-oblivious computation, where communication is reduced to sending and receiving pulses. Algorithms in this setting operate entirely based on the patterns of pulse arrivals.
@@
A natural approach to designing a content-oblivious algorithm is to use unary encoding, representing messages as sequences of pulses. However, in the asynchronous setting, where no upper bound on message delivery time is known, there is no reliable way to detect the end of a sequence. Censor-Hillel, Cohen, Gelles, and Sela [fully-defective-22] established a strong impossibility result: Let $f(x,y)$ be any non-constant function, then any two-party deterministic asynchronous content-oblivious communication protocol, where one party holds $x$ and the other holds $y$, must fail to compute $f$ correctly.
@@
Interestingly, Censor-Hillel, Cohen, Gelles, and Sela [fully-defective-22] showed that in 2-edge-connected networks, any algorithm designed for the noiseless setting can be simulated in a fully defective network. We briefly sketch the main ideas underlying their approach below.
@@
To extend this approach to general 2-edge-connected networks, they leverage a result by Robbins [robbins1939theorem], which states that any 2-edge-connected undirected graph can be oriented to form a strongly connected directed graph. This guarantees the existence of a Robbins cycle, which is an oriented ring that traverses all nodes, possibly revisiting some nodes multiple times. They designed a content-oblivious algorithm that uses a depth-first search (DFS) to construct such a cycle.
@@
A key limitation of their result is that it inherently relies on a preselected leader. For instance, a leader is needed to generate a unique token in a ring or to select the root for the DFS. Censor-Hillel, Cohen, Gelles, and Sela [fully-defective-22] conjectured that the preselected leader assumption is essential for any non-trivial content-oblivious computation, leaving the following question open.
@@
To answer the above question affirmatively, it suffices to design a content-oblivious leader election algorithm that possesses certain desirable properties, enabling it to be composed with other algorithms. Recently, Frei, Gelles, Ghazy, and Nolin [content-oblivious-leader-election-24] answered this question for the special case of oriented rings. They designed two asynchronous content-oblivious leader election algorithms, both with message complexity $O(n\cdot\mathsf{ID_{\max}})$, where $n$ is the number of nodes and $\mathsf{ID_{\max}}$ is the maximum identifier. The first algorithm stabilizes in unoriented rings, without termination detection. The second algorithm quiescently terminates in oriented rings, ensuring that no further pulses are sent to a node $v$ after the algorithm on $v$ declares termination.
The quiescent termination property enables the execution of the general simulation algorithm discussed earlier, following their leader election algorithm, as each node can correctly identify which pulses belong to which algorithm. As a result, this finding refutes the conjecture of Censor-Hillel, Cohen, Gelles, and Sela [fully-defective-22] for the special case of oriented ring topologies. Given the work by Frei, Gelles, Ghazy, and Nolin [content-oblivious-leader-election-24], the next question to address is whether content-oblivious leader election is feasible for general 2-edge-connected networks.
@@
Existing approaches for content-oblivious leader election heavily rely on the ring structure, making use of its two orientations to handle two types of messages. We briefly outline the key ideas behind the algorithms of Frei, Gelles, Ghazy, and Nolin [content-oblivious-leader-election-24] as follows.
@@
1.1 Our Contribution
We answer 1 affirmatively under the assumption that an upper bound $N\geq n$ on the number of nodes $n=|V|$ is known, by presenting a content-oblivious leader election algorithm that quiescently terminates on any 2-edge-connected network.
@@ 1.1 Our Contribution
Combined with the simulation result of Censor-Hillel, Cohen, Gelles, and Sela [fully-defective-22], our finding implies [^4] More precisely, as established in the prior work [content-oblivious-leader-election-24], a sufficient condition for achieving this objective consists of (1) quiescent termination and (2) the requirement that the leader is the final node to terminate. This is because the algorithmic simulation in [fully-defective-22] is initiated by the leader. that in any 2-edge-connected network, any algorithm designed for the noiseless setting can be simulated in the fully-defective setting, without assuming a preselected leader, answering 2 affirmatively, again under the assumption that an upper bound $N\geq n$ is known. This refutes the original conjecture posed by Censor-Hillel, Cohen, Gelles, and Sela [fully-defective-22].
@@
Unoriented rings Stabilizing $O\left(n\cdot\mathsf{ID_{\max}}\right)$$\times$ [content-oblivious-leader-election-24]
Oriented rings Quiescently terminating $O\left(n\cdot\mathsf{ID_{\max}}\right)$$\times$ [content-oblivious-leader-election-24]
Oriented rings Stabilizing $\Omega\left(n\cdot\log\frac{\mathsf{ID_{\max}}}{n}\right)$$\times$ [content-oblivious-leader-election-24]
2-edge-connected Quiescently terminating $O\left(m\cdot N\cdot\mathsf{ID_{\min}}\right)$ Known $N$ Theorem 1
See Table 1 for a comparison between Theorem 1 and the results from prior work [content-oblivious-leader-election-24]. Although our algorithm has a higher message complexity of $O(m\cdot N\cdot\mathsf{ID_{\min}})$ compared to the previous $O(n\cdot\mathsf{ID_{\max}})$ bound in terms of network size, it reduces the dependency on identifiers from $\mathsf{ID_{\max}}$ to $\mathsf{ID_{\min}}$.
At first glance, our upper bound of $O(m\cdot N\cdot\mathsf{ID_{\min}})$ may seem to contradict the lower bound of $\Omega\left(n\cdot\log\frac{\mathsf{ID_{\max}}}{n}\right)$ established in previous work [content-oblivious-leader-election-24]. However, the two results are consistent, as the lower bound does not make any assumptions about the value of $\mathsf{ID_{\min}}$. Precisely, their bound is of the form $\Omega\left(n\cdot\log\frac{k}{n}\right)$, where $k$ is the number of distinct, assignable identifiers in the network. In other words, $k$ is the size of the $\mathsf{ID}$ space.
While our algorithm offers improvements, it does require a mild assumption: All nodes must a priori agree on a known upper bound $N$ on the number of nodes $n=|V|$ in the network, making it non-uniform. This assumption is widely adopted in many existing leader election and distributed graph algorithms, as upper bounds on network size are often available or can be estimated. In contrast, the algorithms from prior work [content-oblivious-leader-election-24] are uniform, requiring no prior knowledge of the network.
@@ 1.2 Additional Related Work
To address message corruption, a common approach is to introduce additional redundancy using coding techniques, known as interactive coding. The noise affecting the messages can be either random or adversarial, and there is usually an upper bound on the level of randomness or adversarial behavior. The study of interactive coding was initiated by Schulman [schulman1992communication, schulman1993deterministic, 556671] in the two-party setting and later extended to the multi-party case by Rajagopalan and Schulman [rajagopalan1994coding]. For a comprehensive overview of interactive coding, see the survey by Gelles [gelles2017coding].
@@
For networks with $n$ nodes, the maximum fraction of corrupted messages that a distributed protocol can tolerate is $\Theta(1/n)$ [jain2015interactive]. If more than this fraction of messages are corrupted, the adversary can fully disrupt the communication of the node that communicates the fewest messages. Censor-Hillel, Gelles, and Haeupler [censor2019making] presented a distributed interactive coding scheme that simulates any asynchronous distributed protocol while tolerating an optimal corruption of $\Theta(1/n)$ of all messages. A key technique underlying their algorithm is a content-oblivious BFS algorithm.
@@
The celebrated impossibility theorem of Fischer, Lynch, and Paterson [fischer1985impossibility] states that achieving consensus in an asynchronous distributed system is impossible for a deterministic algorithm when one or more nodes may crash. If the number of Byzantine edges is bounded by $f$, reliable communication can only be achieved if the graph is at least $2f$-edge-connected [dolev1982byzantine, pelc1992reliable].
@@
A series of recent studies has focused on resilient and secure distributed graph algorithms in the synchronous setting [DBLP:conf/podc/ParterY19, DBLP:conf/soda/ParterY19a, DBLP:conf/wdag/HitronP21a, DBLP:conf/wdag/HitronP21, DBLP:conf/wdag/HitronPY22, DBLP:conf/podc/Parter22, DBLP:conf/podc/FischerP23, DBLP:conf/innovations/HitronPY23]. These works have developed compilation schemes that transform standard distributed algorithms into resilient and secure versions. For instance, in a $(2f+1)$-edge-connected network, any distributed algorithm in the $\mathsf{CONGEST}$ model can be adapted to remain resilient against up to $f$ adversarial edges [DBLP:conf/wdag/HitronP21a].
@@
Content-oblivious computation has also been studied in the synchronous setting. In the beeping model introduced by Cornejo and Kuhn [beeping-model-10], during each communication round, a node can either beep or listen. If the node listens, it can distinguish between silence or the presence of at least one beep. Several leader election algorithms have been developed for the beeping model and its variations [beeping-multi-hop-radio-networks-19, beeping-time-optimal-18, forster2014deterministic, ghaffari2013near, beeping-minimal-weak-communication-25].
@@
Several additional distributed models have been designed with limited communication capabilities to capture various real-world constraints. These include radio networks [chlamtac2003broadcasting], population protocols [angluin2006computation], and stone-age distributed computing [emek2013stone]. Leader election has been extensively studied in these models [alistarh2022near, berenbrink2020optimal, beeping-multi-hop-radio-networks-19, czumaj2021exploiting, ChangKPWZ17, doty2018stable, emek2021thin, ghaffari2013near, sudo2020leader, beeping-minimal-weak-communication-25].
@@ 2.1 Content-Oblivious Model
The communication network is abstracted as a graph $G=(V,E)$, where each node $v\in V$ is a computing device and each edge $e\in E$ is a bidirectional communication link. We write $n=|V|$ and $m=|E|$. Each node $v$ has a unique identifier $\mathsf{ID}(v)$, and we define $\mathsf{ID_{\max}}=\max_{v\in V}\mathsf{ID}(v)$ and $\mathsf{ID_{\min}}=\min_{v\in V}\mathsf{ID}(v)$. Throughout the paper, we assume that $G$ is 2-edge-connected, meaning that the removal of any single edge does not disconnect the graph. We assume no prior knowledge of the network topology, except that each node is given an upper bound $N\geq n$ on the total number of nodes.
The degree of a node $v$, denoted by $\deg(v)$, is the number of edges incident to $v$. Although a node does not initially know the identifiers of its neighbors, it can distinguish among its incident edges using port numbering—a bijection between the set of incident edges and the set $\{1,2,\ldots,\deg(v)\}$.
@@
Throughout the paper, we write $[s]=\{1,2,\ldots,s\}$. For each node $v\in V$ and each port number $i\in[\deg(v)]$, we use the following notation to track the number of pulses sent and received so far:

### v3

@@ Abstract
In this work, we present two results:
General 2-edge-connected topologies:
First, we show an asynchronous content-oblivious leader election algorithm that quiescently terminates in any 2-edge-connected network with message complexity $O(m\cdot N\cdot\mathsf{ID_{\min}})$, where $m$ is the number of edges, $N$ is a known upper bound on the number of nodes, and $\mathsf{ID_{\min}}$ is the smallest $\mathsf{ID}$.
Combined with the above simulation, this result shows that whenever a size bound $N$ is known, any noiseless algorithm can be simulated in the fully-defective model without a preselected leader, fully refuting the conjecture.
Unoriented rings:
We then show that the knowledge of $N$ can be dropped in unoriented ring topologies by presenting a quiescently terminating election algorithm with message complexity $O(n\cdot\mathsf{ID}_{\max})$ that matches the previous bound.
Consequently, this result constitutes a strict improvement over the previous state of the art and shows that, on rings, fully-defective and noiseless communication are computationally equivalent, with no additional assumptions.
@@ 1 Introduction
Fault tolerance is a cornerstone of distributed computing, enabling systems to remain operational despite various failures, such as node crashes or channel noise [BDM93, Dub13, KK20, Ray18]. One important category of faults, known as alteration errors, changes the content of messages but does not create new messages or eliminate existing ones. Traditional methods for handling such errors often rely on adding redundancy through coding techniques and assume known bounds on the frequency or severity of faults. However, in many practical settings, such assumptions may not hold, limiting the effectiveness of these approaches.
@@
Censor-Hillel, Cohen, Gelles, and Sela [CCG+23] introduced the model of fully-defective asynchronous networks, in which all links are subject to a severe form of alteration errors: The content of every message can be arbitrarily modified, although the adversary cannot insert new messages or remove existing messages. Since messages can no longer carry meaningful information, the model is equivalent to content-oblivious computation, where communication is reduced to sending and receiving pulses. Algorithms in this setting operate entirely based on the patterns of pulse arrivals.
@@
A natural approach to designing a content-oblivious algorithm is to use unary encoding, representing messages as sequences of pulses. However, in the asynchronous setting, where no upper bound on message delivery time is known, there is no reliable way to detect the end of a sequence. Censor-Hillel, Cohen, Gelles, and Sela [CCG+23] established a strong impossibility result: Let $f(x,y)$ be any non-constant function, then any two-party deterministic asynchronous content-oblivious communication protocol, where one party holds $x$ and the other holds $y$, must fail to compute $f$ correctly.
@@
Interestingly, Censor-Hillel, Cohen, Gelles, and Sela [CCG+23] showed that in 2-edge-connected networks, any algorithm designed for the noiseless setting can be simulated in a fully defective network. We briefly sketch the main ideas underlying their approach below.
@@
To extend this approach to general 2-edge-connected networks, they leverage a result by Robbins [Rob39], which states that any 2-edge-connected undirected graph can be oriented to form a strongly connected directed graph. This guarantees the existence of a Robbins cycle, which is an oriented ring that traverses all nodes, possibly revisiting some nodes multiple times. They designed a content-oblivious algorithm that uses a depth-first search (DFS) to construct such a cycle.
@@
A key limitation of their result is that it inherently relies on a preselected leader. For instance, a leader is needed to generate a unique token in a ring or to select the root for the DFS. Censor-Hillel, Cohen, Gelles, and Sela [CCG+23] conjectured that the preselected leader assumption is essential for any non-trivial content-oblivious computation, leaving the following question open.
@@
To answer the above question affirmatively, it suffices to design a content-oblivious leader election algorithm that possesses certain desirable properties, enabling it to be composed with other algorithms. Recently, Frei, Gelles, Ghazy, and Nolin [FGG+24] answered this question for the special case of oriented rings. They designed two asynchronous content-oblivious leader election algorithms, both with message complexity $O(n\cdot\mathsf{ID_{\max}})$, where $n$ is the number of nodes and $\mathsf{ID_{\max}}$ is the maximum identifier. The first algorithm stabilizes in unoriented rings, without termination detection. The second algorithm quiescently terminates in oriented rings, ensuring that no further pulses are sent to a node $v$ after the algorithm on $v$ declares termination.
The quiescent termination property enables the execution of the general simulation algorithm discussed earlier, following their leader election algorithm, as each node can correctly identify which pulses belong to which algorithm. As a result, this finding refutes the conjecture of Censor-Hillel, Cohen, Gelles, and Sela [CCG+23] for the special case of oriented ring topologies. Given the work by Frei, Gelles, Ghazy, and Nolin [FGG+24], the next question to address is whether content-oblivious leader election is feasible for general 2-edge-connected networks.
@@
Existing approaches for content-oblivious leader election heavily rely on the oriented ring structure, making use of its two orientations to handle two types of messages. We briefly outline the key ideas behind the algorithms of Frei, Gelles, Ghazy, and Nolin [FGG+24] as follows.
@@
We stress that orientation is a key assumption of the above algorithm. In fact, they conjectured the impossibility of terminating leader election in unoriented rings [FGG+24].
@@
1.1 Contributions
We answer ˜1 affirmatively under the assumption that an upper bound $N\geq n$ on the number of nodes $n=|V|$ is known, by presenting a content-oblivious leader election algorithm that quiescently terminates on any 2-edge-connected network.
@@ 1.1 Contributions
Combined with the simulation result of Censor-Hillel, Cohen, Gelles, and Sela [CCG+23], our finding implies that in any 2-edge-connected network, any algorithm designed for the noiseless setting can be simulated in the fully-defective setting, without assuming a preselected leader. More precisely, as established in the prior work [FGG+24], a sufficient condition for achieving this objective consists of (1) quiescent termination and (2) the requirement that the leader is the final node to terminate. This is because the algorithmic simulation in [CCG+23] is initiated by the leader. Hence we answer ˜2 affirmatively, again under the assumption that an upper bound $N\geq n$ is known. This refutes the original conjecture posed by Censor-Hillel, Cohen, Gelles, and Sela [CCG+23].
For the special case of ring topologies, we can drop the assumption of knowing $N$ without requiring orientation.
Theorem 2.
There exists a quiescently terminating leader election algorithm with message complexity $O(n\cdot\mathsf{ID_{\max}})$ in any unoriented ring, where $\mathsf{ID_{\max}}$ is the largest identifier. Moreover, the leader is the last node to terminate.
This result refutes the conjecture of [FGG+24], demonstrating that orientation is not necessary for quiescently terminating leader election in rings.
@@
Unoriented rings Stabilizing $O\left(n\cdot\mathsf{ID_{\max}}\right)$ $\times$ [FGG+24]
Oriented rings Quiescently terminating $O\left(n\cdot\mathsf{ID_{\max}}\right)$ $\times$ [FGG+24]
Oriented rings Stabilizing $\Omega\left(n\cdot\log\frac{\mathsf{ID_{\max}}}{n}\right)$ $\times$ [FGG+24]
2-edge-connected Quiescently terminating $O\left(m\cdot N\cdot\mathsf{ID_{\min}}\right)$ Known NN Theorem˜1
Unoriented rings Quiescently terminating $O\left(n\cdot\mathsf{ID_{\max}}\right)$ $\times$ Theorem˜2
See Table˜1 for a comparison of our results and the results from prior work [FGG+24].
Our algorithm for general 2-edge-connected topologies has a higher message complexity of $O(m\cdot N\cdot\mathsf{ID_{\min}})$ compared to the previous $O(n\cdot\mathsf{ID_{\max}})$ bound in terms of network size, but it reduces the dependency on identifiers from $\mathsf{ID_{\max}}$ to $\mathsf{ID_{\min}}$.
At first glance, the upper bound of $O(m\cdot N\cdot\mathsf{ID_{\min}})$ may seem to contradict the lower bound of $\Omega\left(n\cdot\log\frac{\mathsf{ID_{\max}}}{n}\right)$ established in previous work [FGG+24].
However, the two results are consistent, as the lower bound does not make any assumptions about the value of $\mathsf{ID_{\min}}$. Precisely, their bound is of the form $\Omega\left(n\cdot\log\frac{k}{n}\right)$, where $k$ is the number of distinct, assignable identifiers in the network. In other words, $k$ is the size of the $\mathsf{ID}$ space.
While our algorithm for Theorem˜1 solves leader election in any 2-edge-connected topology,
it does require a mild assumption: All nodes must a priori agree on a known upper bound $N$ on the number of nodes $n=|V|$ in the network, making it non-uniform. This assumption is widely adopted in many existing leader election and distributed graph algorithms, as upper bounds on network size are often available or can be estimated. In contrast, the terminating algorithm from prior work [FGG+24] is uniform, requiring no prior knowledge of the network, but it only works on oriented rings.
Our leader election algorithm on rings requires neither orientation nor knowledge of $N$, while matching the message complexity of the algorithms of [FGG+24]. It therefore constitutes a strict improvement over the previous state of the art.
@@ 1.2 Additional Related Work
To address message corruption, a common approach is to introduce additional redundancy using coding techniques, known as interactive coding. The noise affecting the messages can be either random or adversarial, and there is usually an upper bound on the level of randomness or adversarial behavior. The study of interactive coding was initiated by Schulman [Sch92, Sch93, Sch96] in the two-party setting and later extended to the multi-party case by Rajagopalan and Schulman [RS94]. For a comprehensive overview of interactive coding, see the survey by Gelles [Gel17].
@@
For networks with $n$ nodes, the maximum fraction of corrupted messages that a distributed protocol can tolerate is $\Theta(1/n)$ [JKL15]. If more than this fraction of messages are corrupted, the adversary can fully disrupt the communication of the node that communicates the fewest messages. Censor-Hillel, Gelles, and Haeupler [CGH19] presented a distributed interactive coding scheme that simulates any asynchronous distributed protocol while tolerating an optimal corruption of $\Theta(1/n)$ of all messages. A key technique underlying their algorithm is a content-oblivious BFS algorithm.
@@
The celebrated impossibility theorem of Fischer, Lynch, and Paterson [FLP85] states that achieving consensus in an asynchronous distributed system is impossible for a deterministic algorithm when one or more nodes may crash. Similar impossibility results on the solvability of consensus have been found when processes are correct but may experience communication failures, such as message omissions, insertions, and corruptions [SW89].
If the number of Byzantine edges is bounded by $f$, reliable communication can only be achieved if the graph is at least $2f$-edge-connected [Dol82, Pel92].
@@
A series of recent studies has focused on resilient and secure distributed graph algorithms in the synchronous setting [PY19a, PY19, HP21a, HP21, HPY22, Par22, FP23, HPY23]. These works have developed compilation schemes that transform standard distributed algorithms into resilient and secure versions. For instance, in a $(2f+1)$-edge-connected network, any distributed algorithm in the $\mathsf{CONGEST}$ model can be adapted to remain resilient against up to $f$ adversarial edges [HP21a].
@@
Content-oblivious computation has also been studied in the synchronous setting. In the beeping model introduced by Cornejo and Kuhn [CK10], during each communication round, a node can either beep or listen. If the node listens, it can distinguish between silence or the presence of at least one beep. Several leader election algorithms have been developed for the beeping model and its variations [CD19, DBB18, FSW14, GH13, VZ25].
@@
Several additional distributed models have been designed with limited communication capabilities to capture various real-world constraints. These include radio networks [CK03], population protocols [AAD+06], and stone-age distributed computing [EW13]. Leader election has been extensively studied in these models [ARV22, BGK20, CD19, CD21, CKP+19, DS18, EK21, GH13, SM20, VZ25].
@@ 2.1 Content-Oblivious Model
The communication network is abstracted as a graph $G=(V,E)$, where each node $v\in V$ is a computing device and each edge $e\in E$ is a bidirectional communication link. We write $n=|V|$ and $m=|E|$. Each node $v$ has a unique identifier $\mathsf{ID}(v)$, and we define $\mathsf{ID_{\max}}=\max_{v\in V}\mathsf{ID}(v)$ and $\mathsf{ID_{\min}}=\min_{v\in V}\mathsf{ID}(v)$. Throughout the paper, we assume that $G$ is 2-edge-connected, meaning that the removal of any single edge does not disconnect the graph.
General 2-edge-connected topologies.
For our result on general 2-edge-connected topologies, we assume no prior knowledge of the network topology, except that each node is given an upper bound $N\geq n$ on the total number of nodes. That is, the algorithm is non-uniform.
The degree of a node $v$, denoted by $\deg(v)$, is the number of edges incident to $v$. Although a node does not initially know the identifiers of its neighbors, it can distinguish among its incident edges using port numbering—a bijection between the set of incident edges and the set $\{1,2,\ldots,\deg(v)\}$. Therefore, we use the term port to refer to the local endpoint of an edge, uniquely identified by its port number, which is used to send and receive messages to and from a specific neighbour.
Unoriented rings.
For our result on rings, we assume that the topology is an unoriented ring. In this case, each node $v_{i}$ is connected to nodes $v_{i-1}$ and $v_{i+1}$ (indices taken modulo $n$) by two local communication ports: port 0 and port $1$. These local labels do not induce a consistent orientation of the ring; that is, the global assignment of labels 0 and $1$ to ports is arbitrarily set at the beginning by an adversary. It may be the case that a message sent by $v_{1}$ on port 0 is delivered to node $v_{0}$ on its local port $1$, and a message sent by $v_{1}$ on port $1$ is delivered to $v_{2}$ on its local port $1$. A correct algorithm must work for any possible assignment of port labels.
Content-oblivious communication.
@@
We write $s]=\{1,2,\ldots,s\}$. In the algorithm description and analysis for Theorem˜1, for each node $v\in V$ and each port number $i\in[\deg(v)]$, we use the following notation to track the number of pulses sent and received so far: